OpenStudy (toneyaarionnepeyton):

You toss 5 coins. What is the probability that all 5 show heads? 1/8 1/32 1/16 1/12 1/4

OpenStudy (cas_fangirl_14):

Each coin has a 1 in 2 chance of landing heads so probability for each one is=1/2 You have tosses 5 coins so for 5 coins it will be =(1/2)*(1/2)*(1/2)*(1/2)*(1/2)=1/32

OpenStudy (toneyaarionnepeyton):

omg thank you lol

OpenStudy (cas_fangirl_14):

no problem :)

OpenStudy (cas_fangirl_14):

You should close the question unless you need more help.